2x2 – 5x + 7 = 0 (1 point)
31; one, real
31; two, real
–31; one, imaginary
–31; two, imaginary
6. –3x2 + 14x – 8 = 0 (1 point)
100; one, real
100; two, real
196; one, real
196; two, real
7. 4x2 – 5x + 10 = 7x + 1 (1 point)
0; one, real
0; no real solutions
–135, one, imaginary
–135; two, imaginary

Answer by Alan3354(69443)

2x2 – 5x + 7 = 0 (1 point)
31; one, real
31; two, real
–31; one, imaginary
–31; two, imaginary
6. –3x2 + 14x – 8 = 0 (1 point)
100; one, real
100; two, real
196; one, real
196; two, real
7. 4x2 – 5x + 10 = 7x + 1 (1 point)
0; one, real
0; no real solutions
–135, one, imaginary
–135; two, imaginary
===================
What's your question?